OpenStudy (anonymous):

find the equation of the line useing the following info show all your work (-7,-7)(-9,9)

OpenStudy (anonymous):

Start by finding the slope using the slope formula.

OpenStudy (anonymous):

(y-y1)=((y2-y1)/(x2-x1)) (x-x1) y+7=((9+7/-9+7))(x+7) y+7=(16/-2)(x+7) y+7=-8(x+7) y+7=-8x-56 8x+y=-56-7 8x+y=-63